OpenStudy (anonymous):

Devin won a tiny bouncy ball at the school carnival. The diameter of the ball is 1.25 inches. To the nearest hundredth of a square inch, what is the surface area of the ball? Use 3.14 for pie

OpenStudy (callisto):

assume it is a perfect sphere surface area = 4(pi)r^2 = 4(pi)(diamter/2)^2, can you substitute the number and work out the answer yourself?